OpenStudy (anonymous):

One More Log Question Logbase5(2x+5)=2

OpenStudy (anonymous):

So we take 5 and raise it to the 2nd power, why is that? I just can't wrap my head around it. Maybe because i am picturing it like this

OpenStudy (zzr0ck3r):

raise both side to be powers of 5 5^(log_5(2x+5) = 5^2 2x+5 = 25

OpenStudy (zzr0ck3r):

you have a = b I did 5^a = 5^b

OpenStudy (anonymous):

\[\log_{5} 2\] This does not equal 25, am I just picturing it wrong?

OpenStudy (zzr0ck3r):

no it does not

OpenStudy (zzr0ck3r):

5^? = 2

OpenStudy (zzr0ck3r):

you are not taking the log of both sides, you are raising them both to the base of the log

OpenStudy (anonymous):

Okay so i am using a log cancellation property to get rid of the log on the left hand side?

OpenStudy (anonymous):

x=e^1 then right?

OpenStudy (zzr0ck3r):

yeah lhs 5^(log_5(2x+5)) = 2x+5

OpenStudy (zzr0ck3r):

yeah sorry :)

OpenStudy (anonymous):

:) awesome man, so i raise it to the base to cancel out the log, then solve it like i normally would?

OpenStudy (zzr0ck3r):

yeah